OpenStudy (anonymous):

The graphs below show four functions:

OpenStudy (igreen):

Please post the whole question..

OpenStudy (anonymous):

Which graph best shows the function f(x) = 3(3)-x to represent the rate at which a radioactive substance decays? (A) Graph A (B) Graph B (C) Graph C (D) Graph D

OpenStudy (anonymous):

Is this C? @ganeshie8 @hartnn @ParthKohli @iambatman

hartnn (hartnn):

if it is \(\Large 3^{3-x}\) then no

OpenStudy (anonymous):

Oh, ok. Well, a graph that shows decay is negative, right?

OpenStudy (igreen):

^^^

OpenStudy (igreen):

Correct. Decay = Negative

OpenStudy (igreen):

So which one's can we take out?

hartnn (hartnn):

is it \( 3^{3-x} \) ?

OpenStudy (anonymous):

Srry this is actually 3(3)^(-x)

OpenStudy (igreen):

@hartnn I think it's \(f(x) = 3(3)^{-x}\).

OpenStudy (igreen):

Which lines can you take out since you know that the line must be negative?

OpenStudy (anonymous):

\[3(3)^{-x}\]

hartnn (hartnn):

ok, then C is correct :)

OpenStudy (igreen):

Yes.. We can take out A & B, since they are positive.. \(f(x) = 3(3)^{-x}\) The 3 before the parenthesis tells us the y-intercept..which is 3..and line C has a y-intercept of 3..so it must be our answer.

OpenStudy (anonymous):

We can take out A and B

OpenStudy (igreen):

And D has a y-intercept of 1..so it can't be our answer.

OpenStudy (anonymous):

Yup that's what I said first. My answer was C, but the way I typed the expression was incorrect. Then I typed it the correct way. :)

OpenStudy (igreen):

:)